Simplify the following expressions:
(a). (cos π/4 + isin π/4)(cos 3π/4 + isin 3π/4), (b). (cos π/4 + isin π/4)2
(cos π/6 + isin π/6) .
1
Expert's answer
2020-05-20T19:59:32-0400


(a). (cos π/4 + isin π/4)(cos 3π/4 + isin 3π/4)=( "\sqrt{2}"/2+i"\sqrt{2}"/2)(- "\sqrt{2}"/2+i"\sqrt{2}"/2)= -2/4+i*2/4-i*2/4+i2*2/4=-1/2+i2*1/2= -1/2 -1/2=-2/2=-1

(b). (cos π/4 + isin π/4)2(cos π/6 + isin π/6)= ( "\sqrt{2}"/2+ i"\sqrt{2}"/2)*2*("\sqrt{3}"/2 + i/2)=("\sqrt{2}"/2+ i"\sqrt{2}"/2)("\sqrt{3}"+i)= "\sqrt{6}"/2 + i*"\sqrt{2}"/2 + i*"\sqrt{6}"/2 + i2*"\sqrt{2}"/2="\sqrt{6}"/2 - "\sqrt{2}"/2 + i*"\sqrt{2}"/2 + i*"\sqrt{6}"/2= ("\sqrt{2}"("\sqrt{3}" -1) + "\sqrt{2}"(1+"\sqrt{3}" ))/2= ("\sqrt{2}" (3-1))/2= 2"\sqrt{2}"/2 = "\sqrt{2}"

Answer: (a). (cos π/4 + isin π/4)(cos 3π/4 + isin 3π/4)=-1, (b). (cos π/4 + isin π/4)2(cos π/6 + isin π/6)="\sqrt{2}"